# SASG-DA: Sparse-Aware Semantic-Guided Diffusion Augmentation For Myoelectric Gesture Recognition

### v2

@@ 5.1 Datasets
Baseline 74.63 75.89 75.44 75.66 70.00 71.44 70.83 71.13 69.71 71.86 70.83 71.34
Jittering & Scaling 74.36 75.75 75.20 75.47 70.08 71.67 70.94 71.30 68.46 70.78 69.61 70.19
Upsample 75.84 77.18 76.85 77.01 70.05 71.66 70.91 71.28 74.38 76.69 75.36 76.02
FreqMask 74.37 75.90 75.36 75.63 71.22 72.90 72.16 72.53 73.86 75.70 74.89 75.29
RIM 75.69 76.81 76.53 76.67 71.36 72.72 72.28 72.50 72.14 73.78 73.20 73.49
Dominant Shuffle 75.82 77.10 76.63 76.86 71.02 73.01 71.91 72.46 74.38 76.17 75.33 75.75
Mixup 74.45 75.65 75.45 75.55 71.13 72.47 72.12 72.29 69.46 71.44 70.58 71.01
STAug 74.49 75.65 75.40 75.52 71.02 72.36 71.94 72.15 69.52 71.45 70.72 71.08
FreqMix 77.07 78.57 77.86 78.21 71.47 73.56 72.33 72.94 73.92 75.97 74.94 75.45
SimPSI 74.58 75.74 75.42 75.58 70.52 71.98 71.43 71.70 70.91 72.80 72.00 72.40
DiffMix 74.84 76.07 75.74 75.90 69.47 70.86 70.49 70.67 70.33 72.01 71.49 71.75
Minority 76.89 77.59 77.78 77.68 71.31 72.34 72.16 72.25 73.48 74.57 74.52 74.54
CADS 77.03 77.74 77.87 77.80 72.33 73.27 73.08 73.17 74.33 75.34 75.29 75.31
SASG-DA (Ours) 77.86 78.82 78.73 78.77 74.31 75.46 75.20 75.33 76.36 77.59 77.29 77.44
@@ 5.2 Evaluation Metrics
We evaluate the classification performance under the intra-subject setting, where each subject is trained and tested separately. Accuracy (ACC) is reported as the average classification accuracy across all subjects and serves as the primary indicator of the model’s classification ability. In addition, the average precision (Pre), recall (Rec), and F1-score are calculated across all subjects to offer a more detailed assessment of the model’s classification performance.
@@
To assess the quality of samples generated by the diffusion model, we adopt the Fréchet Inception Distance (FID) [heusel2017gans] and Category Accuracy Score (CAS). FID quantifies the distributional similarity between real and generated samples in a latent feature space. CAS evaluates the semantic consistency of generated samples by measuring the classification accuracy of generated data using a classifier trained solely on real samples. It reflects the model’s ability to generate the intended target classes.
@@
To further analyze the diversity and sparsity of generated samples, we employ three sparsity-aware metrics derived from [sehwag2022generating, um2024self], following the same evaluation settings: (1) Average KNN Distance (AvgKNN) measures the average distance from a sample to its nearest neighbors in the feature space ($k=5$). (2) Local Outlier Factor (LOF) [breunig2000lof] assesses how the local density of a sample deviates from that of its neighbors ($k=20$). (3) Rarity Score [han2024rarity] explicitly quantifies the rarity of individual generated samples ($k=5$). We compare the sparsity of generated samples against that of real samples.
5.3 Baseline Models
@@
To evaluate the data augmentation performance of our proposed approach, we adopt three backbone models: Crossformer, TDCT, and STCNet.
Crossformer [zhang2023crossformer] is a popular Transformer-based model for time series forecasting, which has also demonstrated superior performance in sEMG classification tasks due to its ability to effectively model cross-time and cross-channel dependencies and extract multi-scale temporal features. In our experiments, we set the segment length to 16, the window size to 2, and the number of layers to 5.
TDCT [wang2024transformer] is a Transformer-based model tailored for sEMG-based gesture recognition. It incorporates temporal depthwise convolution into the self-attention mechanism to enhance sequence modeling while reducing parameters and computational cost. We follow the original architecture and settings from [wang2024transformer].
STCNet [yang2025stcnet] is also specifically developed for sEMG-based gesture recognition. It combines convolutional and recurrent layers with a spatio-temporal block to enhance feature extraction across time and space. It also incorporates a rolling convolution module to capture the spatial structure inherent in the placement of the sEMG sensor. The input is reshaped to $(C,20,20)$ following [yang2025stcnet], and other configurations remain unchanged.
@@
Augmentation Label condition GMSS SASS DB7 DB4
@@
5.4 Experimental Settings
The experiments are divided into two parts: Diffusion-based data augmentation and downstream classification. The diffusion model is implemented based on a 1D U-Net and is trained exclusively on the original training set. It employs 1000 diffusion steps with a cosine noise schedule and is optimized using the $x_{0}$-prediction objective with the Adam optimizer (learning rate: 1e-5, batch size: 128) for 20,000 iterations. A conditional dropout rate of 0.05 is applied during training to enhance generalization. During sampling, we use DDIM with 500 steps to generate synthetic samples. All pretrained classifiers used during the generation process are implemented using Crossformer and are trained solely on the original training set. For sparse-aware sampling, the hyperparameters involved are empirically set to 200 iterations (iter), a neighborhood radius ($\epsilon$) of 3.0, and a learning rate ($\eta$ ) of 10.0 during optimization. The confidence threshold for filtering sparse samples is set to 0.15.
Following the settings in [wang2024enhance], we randomly shuffle the original training data and the generated samples during the experiments. The resulting augmented dataset, which is twice the size of the original training set, is then used for downstream classification. We evaluate the performance using three different backbone classifiers. During training, we employ the SGD optimizer with an initial learning rate of 0.01, which is reduced by a factor of 0.1 at the 60th epoch. The batch size is set to 256, and all models are trained for 100 epochs. All results are reported as the mean over three runs with different random seeds.
5.5 Comparison with the State-of-the-arts
We compare our method against a comprehensive set of state-of-the-art data augmentation approaches. Specifically, we evaluate five single-sample DA methods tailored for time series: Jittering & Scaling [um2017data], Upsample [semenoglou2023data], FreqMask [chen2023fraug], RIM [aboussalah2023recursive], and Dominant Shuffle [zhao2024dominant]. We also consider four mix-based DA methods: Mixup [zhang2018mixup], STAug [zhang2023towards], FreqMix [chen2023fraug], and SimPSI [ryu2024simpsi]. In addition, three diffusion-based DA methods are included. DiffMix [wang2024enhance] is a simple yet effective generative data augmentation strategy that leverages inter-class image interpolation to produce both faithful and diverse samples. Minority [um2024self] leverages gradient guidance to synthesize samples in low-density regions. CADS [sadat2024cads] enhances sample diversity by injecting stochastic perturbations into conditional inputs during the inference process. For single-sample and mix-based DA methods, we adopt a batch-wise augmentation strategy, whereas for diffusion-based methods and RIM, we jointly train the original and augmented samples to preserve data distribution better and improve model generalization. For fairness, all SOTA methods are reproduced using official code when available, or reimplemented based on recommended or optimal settings.
The results in Table 2, Table 3, and Table 4 demonstrate that our method consistently outperforms the state-of-the-art (SOTA) on three datasets and with all three backbones, which confirms its generalizability and consistent effectiveness when used as a data augmentation method. Compared to the next best method, CADS [sadat2024cads], which enhances diversity in diffusion-based augmentation, our method achieves an average accuracy gain of 1.7% across all datasets and backbones. This confirms the effectiveness of guiding sample generation toward sparse regions, resulting in more informative and complementary samples for data augmentation. Compared to Minority [um2024self], our approach achieves better performance by not only improving diversity but also enhancing the faithfulness of generated samples, i.e., their alignment with class-consistent semantics. This improvement in sample fidelity allows the augmented data to better support classifier learning, leading to more substantial gains in accuracy.
Table 6: Hyperparameters ablations on DB7 and DB4 under different settings for three backbones.
iter $\epsilon$ threshold DB7 DB4
@@ 5.6.1 Module Ablation
To evaluate the contribution of each component in our proposed approach, we conduct a series of ablation experiments by progressively integrating each module. The results are shown in Table 5. We first conduct ablation experiments using two baselines: training without any data augmentation and using a simple label-conditioned generation. The results show that data augmentation effectively alleviates the limitations of the training set and improves the performance of gesture recognition. Subsequently, we introduce semantic representations as conditions and adopt the GMSS strategy, which yields significant improvements across all backbones and datasets, demonstrating that the semantic guidance and sampling enhances the faithfulness and diversity of the generated samples. Finally, we incorporate the SASS strategy to explicitly expand the distribution of the training data, which further improves the downstream classification performance. Compared to the baseline, our method achieves an average improvement of 5.8%, confirming its effectiveness and generalizability.
5.6.2 Inter-Subject Variability Ablation
It is well known that sEMG signals exhibit significant subject-level variability due to factors such as muscle anatomy, skin-electrode impedance, and electrode placement. Fig. 4 illustrates the classification accuracy for each subject from the DB7 and DB4 datasets. Our method yields consistent improvements for the majority of subjects, demonstrating its effectiveness in enhancing general performance.
@@ 5.6.3 Hyperparameters Ablation
To better understand the influence of key hyperparameters on the performance of our approach, we conduct ablation studies focusing on three key hyperparameters: the number of iterations (iter) during sparse-aware sampling, the neighborhood radius ($\epsilon$), and the confidence threshold for filtering sparse samples. The first two parameters control the degree of sparsity during sample generation: larger iterations and larger neighborhood radii produce samples with greater sparsity (e.g., higher FID and lower CAS). The confidence threshold serves as a filter mechanism to avoid introducing low-quality samples. The experimental ablations in Table 6 present the results across different settings on DB7 and DB4 using multiple backbone classifiers. We observe that increasing iter and $\epsilon$ slightly affects classification accuracy, suggesting that our sparse-aware sampling strategy is relatively stable under varying sparsity levels. Additionally, performance remains consistent across a range of confidence thresholds, indicating that the filtering mechanism does not overly restrict sample diversity. Overall, these results highlight the robustness of our method concerning key hyperparameter choices and confirm that it generalizes well without requiring sensitive tuning.

### v3

@@ 5.1 Datasets
Baseline $74.63$ p m 6.10∗∗∗ $75.89$ p m 5.68 $75.44$ p m 5.55 $75.66$ p m 5.67 $70.00$ p m 6.19∗∗∗ $71.44$ p m 5.60 $70.83$ p m 5.58 $71.13$ p m 5.65 $69.71$ p m 6.61∗∗∗ $71.86$ p m 5.97 $70.83$ p m 5.93 $71.34$ p m 6.03
Jittering & Scaling $74.36$ p m 6.34∗∗∗ $75.75$ p m 5.84 $75.20$ p m 5.78 $75.47$ p m 5.88 $70.08$ p m 6.34∗∗∗ $71.67$ p m 5.68 $70.94$ p m 5.67 $71.30$ p m 5.75 $68.46$ p m 7.03∗∗∗ $70.78$ p m 6.32 $69.61$ p m 6.30 $70.19$ p m 6.37
[…]
FreqMask $74.37$ p m 6.93∗∗∗ $75.90$ p m 6.27 $75.36$ p m 6.39 $75.63$ p m 6.51 $71.22$ p m 6.01∗∗∗ $72.90$ p m 5.49 $72.16$ p m 5.48 $72.53$ p m 5.57 $73.86$ p m 6.40∗∗∗ $75.70$ p m 5.88 $74.89$ p m 5.89 $75.29$ p m 6.01
RIM $75.69$ p m 6.14∗∗∗ $76.81$ p m 5.75 $76.53$ p m 5.61 $76.67$ p m 5.75 $71.36$ p m 6.34∗∗∗ $72.72$ p m 5.74 $72.28$ p m 5.73 $72.50$ p m 5.82 $72.14$ p m 6.87∗∗∗ $73.78$ p m 6.21 $73.20$ p m 6.21 $73.49$ p m 6.29
Dominant Shuffle $75.82$ p m 6.10∗∗∗ $77.10$ p m 5.69 $76.63$ p m 5.62 $76.86$ p m 5.73 $71.02$ p m 6.58∗∗∗ $73.01$ p m 5.59 $71.91$ p m 5.91 $72.46$ p m 6.01 $74.38$ p m 6.72∗∗∗ $76.17$ p m 6.03 $75.33$ p m 6.04 $75.75$ p m 6.11
Mixup $74.45$ p m 6.28∗∗∗ $75.65$ p m 5.81 $75.45$ p m 5.70 $75.55$ p m 5.82 $71.13$ p m 5.72∗∗∗ $72.47$ p m 5.19 $72.12$ p m 5.15 $72.29$ p m 5.25 $69.46$ p m 6.84∗∗∗ $71.44$ p m 6.12 $70.58$ p m 6.36 $71.01$ p m 6.42
STAug $74.49$ p m 6.44∗∗∗ $75.65$ p m 5.99 $75.40$ p m 5.87 $75.52$ p m 5.99 $71.02$ p m 5.92∗∗∗ $72.36$ p m 5.47 $71.94$ p m 5.31 $72.15$ p m 5.43 $69.52$ p m 7.11∗∗∗ $71.45$ p m 6.53 $70.72$ p m 6.54 $71.08$ p m 6.60
[…]
SimPSI $74.58$ p m 6.06∗∗∗ $75.74$ p m 5.68 $75.42$ p m 5.54 $75.58$ p m 5.67 $70.52$ p m 5.93∗∗∗ $71.98$ p m 5.41 $71.43$ p m 5.28 $71.70$ p m 5.39 $70.91$ p m 7.09∗∗∗ $72.80$ p m 6.43 $72.00$ p m 6.48 $72.40$ p m 6.56
E-TRGAN $75.17$ p m 5.42∗∗∗ $75.85$ p m 6.33 $75.90$ p m 5.96 $75.87$ p m 6.12 $70.41$ p m 5.28∗∗∗ $72.05$ p m 6.41 $71.90$ p m 5.73 $71.98$ p m 6.25 $70.21$ p m 5.15∗∗∗ $71.70$ p m 5.89 $71.82$ p m 5.07 $71.76$ p m 6.48
DiffMix $74.84$ p m 6.28∗∗∗ $76.07$ p m 5.84 $75.74$ p m 5.71 $75.90$ p m 5.86 $69.47$ p m 6.61∗∗∗ $70.86$ p m 6.08 $70.49$ p m 5.98 $70.67$ p m 6.07 $70.33$ p m 6.55∗∗∗ $72.01$ p m 6.11 $71.49$ p m 5.97 $71.75$ p m 6.10
SSE-LDM $74.73$ p m 6.22∗∗∗ $75.88$ p m 5.83 $75.61$ p m 5.68 $75.74$ p m 5.81 $70.28$ p m 6.45∗∗∗ $71.62$ p m 5.97 $71.26$ p m 5.84 $71.44$ p m 5.95 $69.91$ p m 7.01∗∗∗ $71.62$ p m 6.41 $71.08$ p m 6.41 $71.34$ p m 6.51
EEG-Enhancing $75.10$ p m 5.99∗∗∗ $76.35$ p m 5.51 $75.89$ p m 5.45 $76.12$ p m 5.47 $69.93$ p m 6.33∗∗∗ $71.45$ p m 5.79 $70.84$ p m 5.70 $71.15$ p m 5.74 $70.08$ p m 7.05∗∗∗ $72.11$ p m 6.27 $71.22$ p m 6.32 $71.66$ p m 6.29
NTD $75.76$ p m 5.86∗∗∗ $77.14$ p m 5.42 $76.59$ p m 5.31 $76.86$ p m 5.36 $70.87$ p m 6.83∗∗∗ $72.68$ p m 6.01 $71.77$ p m 6.23 $72.22$ p m 6.11 $72.00$ p m 7.26∗∗∗ $73.94$ p m 6.45 $73.02$ p m 6.56 $73.47$ p m 6.50
PatchEMG $75.91$ p m 5.89∗∗∗ $76.96$ p m 5.51 $76.79$ p m 5.40 $76.87$ p m 5.52 $71.31$ p m 6.09∗∗∗ $72.60$ p m 5.59 $72.21$ p m 5.46 $72.40$ p m 5.59 $73.73$ p m 6.40∗∗∗ $75.21$ p m 5.90 $74.74$ p m 5.75 $74.97$ p m 5.90
DESAM $76.32$ p m 6.25∗∗∗ $77.53$ p m 5.82 $77.16$ p m 5.72 $77.34$ p m 5.84 $70.47$ p m 6.41∗∗∗ $72.21$ p m 5.75 $71.32$ p m 5.76 $71.76$ p m 5.81 $72.55$ p m 7.04∗∗∗ $74.41$ p m 6.39 $73.49$ p m 6.40 $73.94$ p m 6.41
Minority $76.89$ p m 5.70∗∗∗ $77.59$ p m 5.40 $77.78$ p m 5.24 $77.68$ p m 5.38 $71.31$ p m 5.84∗∗∗ $72.34$ p m 5.41 $72.16$ p m 5.25 $72.25$ p m 5.37 $73.48$ p m 6.42∗∗∗ $74.57$ p m 5.89 $74.52$ p m 5.83 $74.54$ p m 5.93
[…]
  • $p<0.05$, ∗∗ $p<0.01$, ∗∗∗ $p<0.001$, $N=40$.
@@ 5.2 Evaluation Metrics
We evaluate the classification performance under the intra-subject setting. Accuracy (ACC) is reported as the average classification accuracy across all subjects and serves as the primary indicator of the model’s classification ability. In addition, precision (Pre), recall (Rec), and F1-score are calculated to offer a more detailed assessment of the model’s classification performance. To evaluate whether our method provides statistically meaningful improvements over existing augmentation strategies, we performed subject-wise Wilcoxon signed-rank tests on average accuracy across three random runs per subject for each backbone and applied Bonferroni correction to account for multiple comparisons.
@@
To assess the quality of the generated samples, we adopt the Fréchet Inception Distance (FID) [32] and Category Accuracy Score (CAS). FID quantifies the distributional similarity between real and generated samples in a latent feature space. CAS evaluates the semantic consistency of generated samples by measuring the classification accuracy of generated data using a classifier trained solely on real samples. It reflects the model’s ability to generate the intended target classes.
@@
To further analyze the diversity and sparsity of generated samples, we employ three sparsity-aware metrics derived from [24, 25], following the same evaluation settings: (1) Average KNN Distance (AvgKNN) measures the average distance from a sample to its nearest neighbors in the feature space ($k=5$). (2) Local Outlier Factor (LOF) assesses how the local density of a sample deviates from that of its neighbors ($k=20$). (3) Rarity score [28] explicitly quantifies the rarity of individual generated samples ($k=5$). We compare the sparsity of generated samples against that of real samples.
5.3 Experimental Settings
The experiments consist of diffusion-based data augmentation and downstream classification. The diffusion model is built on a 1D U-Net [18] and trained only on the original training set with 1000 diffusion steps (cosine noise schedule). It uses the $x_{0}$-prediction objective and Adam (lr=1e-5, batch size=128) for 20,000 iterations. A conditional dropout rate of 0.05 is applied. During sampling, DDIM with 500 steps is used. The involved pretrained classifier (Crossformer) during the generation is trained on the original training set for 100 epochs using SGD (lr=0.01), and 128-dimensional semantic representations are extracted from the last fully connected layer. For SASS, hyperparameters are set to 200 iterations (iter), a neighborhood radius ($\epsilon$) of 3.0, a learning rate ($\eta$ ) of 10.0, and a confidence threshold of 0.15.
Following [7], original training and generated samples are shuffled to form an augmented dataset twice the original size, used only for training while test sets remain fully real. To ensure class balance, the synthetic data followed the same class distribution as the training set. Downstream classifiers are trained with SGD (initial lr=0.01, decayed by 0.1 at epoch 60), batch size 256, for 100 epochs. Results are averaged over three runs with different random seeds. All experiments are conducted in PyTorch 2.3.0 on an NVIDIA RTX 4090 GPU.
@@ 5.3 Experimental Settings
To evaluate the data augmentation performance of our proposed approach, we adopt three backbone models: Crossformer [33], TDCT [13], and STCNet [11]. More details about these backbone architectures and the corresponding parameter settings are provided in the Supplementary Material.
5.4 Comparison with the State-of-the-arts
We compare our method against a comprehensive set of state-of-the-art data augmentation approaches. Specifically, we evaluate five single-sample DA methods tailored for time series: Jittering & Scaling [34], Upsample [35], FreqMask [36], RIM [37], and Dominant Shuffle [14]. We also consider four mix-based DA methods: Mixup [38], STAug [39], FreqMix [36], and SimPSI [40]. In addition, diffusion-based DA methods are included. DiffMix [7] is a simple yet effective generative data augmentation strategy that leverages inter-class interpolation to produce diverse samples. Minority [25] leverages gradient guidance to synthesize samples in low-density regions. CADS [23] enhances sample diversity by injecting stochastic perturbations into conditional inputs during the inference process. PatchEMG [6] adopts a patch-based training strategy, and during inference, it generates complete-length signals while employing classifier-free guidance (CFG) to ensure generation quality. SSE-LDM [18], EEG-Enhancing [19], NTD [20], and DESAM [17] are diffusion-based models for EEG signal generation. Among them, DESAM constructs an imputation-based diffusion process and temporal mixup with real samples to improve generation faithfulness. We also include E-TRGAN [41] for comparison, which applies a Transformer-based GAN for sEMG signal reconstruction. For single-sample and mix-based DA methods, we adopt a batch-wise augmentation strategy, whereas for generative-based methods and RIM, we jointly train the original and augmented samples to preserve data distribution better and improve model generalization. For fairness, all SOTA methods are reproduced using official code when available, or reimplemented based on recommended or optimal settings.
The results in Table 2, Table 3, and Table 4 demonstrate that our method consistently outperforms the state-of-the-art (SOTA) on three datasets and with all three backbones, which confirms its generalizability and consistent effectiveness when used as a data augmentation method. Moreover, at the subject level, our method shows statistically significant improvements ($p<0.05$) over nearly all SOTA approaches, with relatively small standard deviations across three runs (0.02–0.11), indicating robust performance (see in the Supplementary Material). Compared to the next best method, CADS [23], which enhances diversity in diffusion-based augmentation, our method achieves an average accuracy gain of 1.7% across all datasets and backbones. This confirms the effectiveness of guiding sample generation toward sparse regions, resulting in more informative and complementary samples for data augmentation. Compared to Minority [25], our approach achieves better performance by not only improving diversity but also enhancing the faithfulness, i.e., their alignment with class-consistent semantics. This improvement in sample faithfulness allows the augmented data to better support classifier learning, leading to more substantial gains in accuracy.
@@ 5.4 Comparison with the State-of-the-arts
Method DB7 DB4
@@
Table 6: Hyperparameters and Semantic Encoder ablations on DB7 and DB4 under different settings for three backbones.
iter $\epsilon$ k threshold DB7 DB4
@@ 5.5.1 Module Ablation
To evaluate the contribution of each component in our proposed approach, we conduct a series of ablation experiments by progressively integrating each module. The results are shown in Table 5. We first conduct ablation experiments using two baselines: training without data augmentation and using a simple label-conditioned generation. The results show that data augmentation effectively alleviates the limitations of the training set and improves the performance of gesture recognition. Subsequently, we introduce semantic representations as conditions and adopt the GMSS strategy, which yields significant improvements across all backbones and datasets, demonstrating that the semantic guidance and sampling enhance the faithfulness and diversity of the generated samples. Finally, we incorporate the SASS strategy to explicitly expand the distribution of the training data, which further improves the downstream classification performance. Compared to the baseline, our method achieves an average improvement of 5.8%, confirming its effectiveness and generalizability. To further analyze model stability for individual subjects under the intra-subject setting, Fig. 4 illustrates the classification accuracy for each subject from the DB7 and DB4 datasets. Our method yields consistent improvements for the majority of subjects, demonstrating its effectiveness in enhancing general performance.
@@ 5.5.2 Hyperparameters Ablation
To better understand the influence of key hyperparameters on the performance, we conduct ablation studies focusing on four key hyperparameters: the number of iterations (iter), the neighborhood radius ($\epsilon$), the neighbor counts $k$ and the confidence threshold for filtering generated samples. The first three parameters jointly control the sparsity level of generated samples: increasing iter corresponds to deeper optimization toward sparse regions, while enlarging $\epsilon$ and $k$ expands the neighborhood considered during sampling and thereby produces samples with greater sparsity (e.g., higher FID and lower CAS). The confidence threshold serves as a filter mechanism to avoid introducing low-quality samples. The experimental ablations in Table 6 present the results across different settings on DB7 and DB4. We observe that increasing iter, $\epsilon$, and $k$ slightly affects classification accuracy, indicating that downstream performance remains relatively stable across a wide range of sparsity settings. This demonstrates the robustness of SASG-DA to sparsity-controlled generation. Additionally, performance remains consistent across a range of confidence thresholds, indicating that the filtering mechanism does not overly restrict sample diversity. Overall, these results highlight the robustness of our method concerning key hyperparameter choices and confirm that it generalizes well without requiring sensitive tuning.
Based on these observations, the default settings can serve as a robust starting point across different datasets, although the exact optimal values may vary slightly. In practice, the neighborhood radius $\epsilon$ is guided by the latent space scale (e.g., comparable to the average inter-sample distance in the training set), while $k$ and iter are related to dataset size and optimization depth (e.g., moderately larger values facilitate sufficient exploration of sparse regions and stable optimization when introducing more samples). Within reasonable ranges, different combinations of these parameters yield comparable downstream performance.
5.5.3 Semantic Encoder Ablation
We further assess the impact of task-aware classifiers on the SRG mechanism using different backbones (Crossformer, TDCT and STCNet). As shown in Table 6, despite differences in test-set performance, the generative metrics show slight variations, while downstream classification performance remains robust, indicating that SRG is not tightly coupled to a specific backbone architecture or to the absolute strength of the pretrained classifier. This behavior can be attributed to the fact that the semantic representations are modeled from the training data, with all three encoders achieving near-saturated training performance and consistently capturing the class-conditional structure. The diffusion model then effectively leverages these representations to generate relatively stable outputs. Crossformer was used as one representative backbone for the main experiments, serving as an alternative among the tested encoders.